OpenStudy (anonymous):

[4_02]Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (apoorvk):

X-intercept is the value of 'x' when y=0, and Y-intercept is value of 'y' when x=0.

OpenStudy (apoorvk):

@SOUTHERNPRIDE - are you even understanding what I am saying or do you just need the answers???

OpenStudy (blockcolder):

\[\frac{x}{a}-\frac{y}{b}=1\] x-intercept is a, y-intercept is b.

OpenStudy (lgbasallote):

y = mx + b b is y-intercept

OpenStudy (apoorvk):

@AndrewNJ - please don;t post the final answer

OpenStudy (anonymous):

x intercept means y=0, so x-2*0=8,x=8 y intercept means x=0, so 0-2y=8, y=-4

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

thanks